An all equity financed company is contemplating the use of debt. Under the following scenario,

A. how much debt financing should be used, and if so why?

Debt/asset ratio(%)           Cost of equity (%)       Common stock share price ($)    EPS

          0                                       5                         20                                 _____

        10                                      6                         22                                  _____

         30                                      8                         23                                  _____

        50                                     10                        21                                  _____

  1.                                     14                       20                                  _____
Answer #1
EPS in a diff. way
Debt/Asset Ratio Cost of equity(%) Common stock share price Value of Equity=MPS/Ke% Value of Debt EPS=MPS*Ke% V(D)+V€ No.of common stock shares EPS
1 2 3 4=3/2 5=4/(100-Col.1)/Col.1 6=2*3 7=4+5 8=4/3 9=4*2/8
0 5 20 400.00 0 1 400 20 1
10 6 22 366.67 40.74 1.32 407.41 17 1.32
30 8 23 287.50 123.21 1.84 410.71 13 1.84
50 10 21 210.00 210.00 2.1 420.00 10 2.1
60 14 20 142.86 214.29 2.8 357.14 7 2.8
Value of debt is calculated below:
The ratio of debt which maximises the value of firm is 50-50 , where, the firm's value is 420
D/(D+E)=0 D/(D+E)=0.5
D/(D+400)=0 D/(D+210)=0.5
D=0 D=210
D/(D+E)=0.1 D/(D+E)=0.6
D/(D+366.67)=0.1 D/(D+142.86)=0.6
D=40.74 D=214.29
D/(D+E)=0.3
D/(D+287.50)=0.3
D=123.21